OpenStudy (anonymous):

The formula (x+a)(x+b) = x^2+(a+b)x is valid. True or False? I say false, but I'm not sure.... Help??

OpenStudy (anonymous):

false

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Your Welcome (:

OpenStudy (zzr0ck3r):

(x+a)(x+b) = x^2 + ax + bx + ab x^2+(a+b)x = x^2 + ax + bx

OpenStudy (zzr0ck3r):

this works if a and b = 0:)

OpenStudy (anonymous):

Thanks @zzr0ck3r Ill try to remember that :D